OpenStudy (arianna1453):

I will give medal and fan!!! Help with calculus problems! (:

OpenStudy (arianna1453):

@lochana

OpenStudy (jango_in_dtown):

use chain rule

OpenStudy (jango_in_dtown):

\[\frac{ d }{ dx }(\sin 3x^2)=\cos 3x^2. \frac{ d }{ dx}(3x^2)\]

OpenStudy (jango_in_dtown):

\[=6x \cos 3x^2\]

OpenStudy (arianna1453):

I go 6xcos(3x^2)

OpenStudy (arianna1453):

Oh okay cool. So what i got was correct. right?

OpenStudy (jango_in_dtown):

yeah

OpenStudy (arianna1453):

Last one. The hardest to me,

OpenStudy (jango_in_dtown):

its easy.. You use the chain rule

OpenStudy (jango_in_dtown):

\[f'(x)=\frac{ 1 }{ x^3+e ^{4x} }.\frac{ d }{ dx }(x^3+e ^{4x})\]

OpenStudy (lochana):

tricky part is that you need to know \[\frac{d(lnx)}{dx} = \frac{1}{x}\]

OpenStudy (lochana):

exactly @jango_IN_DTOWN is correct

OpenStudy (jango_in_dtown):

\[=\frac{ 3x^2+4e^{4x} }{ x^3+e ^{4x} }\]

OpenStudy (jango_in_dtown):

I suggest you read the theory portion of chain rule..

OpenStudy (jango_in_dtown):

Its very useful.. You can do every problem yourself, I am quite sure, once you master it.:)

OpenStudy (jango_in_dtown):

@arianna1453

OpenStudy (lochana):

yes. practice makes you perfect!

OpenStudy (arianna1453):

Thank you guys. I think I will practice that more. So the answer is going to be the 3x^2+4e^4x over x^3+e4x?

OpenStudy (jango_in_dtown):

yeah

OpenStudy (arianna1453):

What you said above basically. It cannot be simplified more, correct?

OpenStudy (jango_in_dtown):

no.. yoou dont have any common terms

OpenStudy (arianna1453):

So you take the outside and multiply it by the derivative of the inside?

OpenStudy (jango_in_dtown):

do you want to learn about the chain rule?

OpenStudy (arianna1453):

Well i know that you take the outside, so ln and multiply it by the derivative of the inside the whole (x^3+e^4x)

OpenStudy (jango_in_dtown):

Lets do the previous problem

OpenStudy (jango_in_dtown):

f(x)= \[\ln (x^3+e ^{4x})\]

OpenStudy (jango_in_dtown):

Let \[X=x^3+e ^{4x}\]

OpenStudy (arianna1453):

RIght,

OpenStudy (jango_in_dtown):

then f'(x)= \[\frac{ d }{ dx}(\ln X)=\frac{ d }{ dX}(\ln X). \frac{ dX }{ dx }\]

OpenStudy (jango_in_dtown):

\[=\frac{ 1 }{ X }.\frac{ d }{ dx }(x^3+e ^{4x})\]

OpenStudy (arianna1453):

okay and then you plug in (x^3+e^4x) in for x. Im following.

OpenStudy (jango_in_dtown):

\[=\frac{ 1 }{ x^3 +e ^{4x}}.(3x^2+4e ^{4x})\]

OpenStudy (jango_in_dtown):

for X

OpenStudy (arianna1453):

I understand it now!

OpenStudy (jango_in_dtown):

yeah you let the entirely big term as X

OpenStudy (arianna1453):

Okay I have one more problem. I am going to try it myself and if I need help Ill tag you.

OpenStudy (jango_in_dtown):

yeah sure

OpenStudy (arianna1453):

Okay so i did everything for this problem. starting with this. \[(d/dx)(x) +2y (dy/dx) = (1)(y) + (x) (dy/dx)\]

OpenStudy (arianna1453):

@jango_IN_DTOWN

OpenStudy (arianna1453):

I got 1-y/x-2y. So i think i messed up with the 2y part

OpenStudy (jango_in_dtown):

first let me know if you know partial derivatives?

OpenStudy (arianna1453):

Like the xy part, correct?

OpenStudy (jango_in_dtown):

wait see d(x+2y)=dx+2dy and d(xy)=xdy+ydx so from x+2y=xy we have dx+2dy=xdy+ydx or, (2-x)dy=(y-1)dx or. dy/dx=(y-1)/(2-x)

OpenStudy (arianna1453):

so where i had 2y (dy/dx) it was wrong there? it should just be 2(dy/dx)

OpenStudy (arianna1453):

I got the right side correct.

OpenStudy (arianna1453):

Anyways, thank you for your help today (:

OpenStudy (jango_in_dtown):

ok.. bye.:)